# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v1

@@ IV-C Brief Description and Settings of Compared Methods
To evaluate the effectiveness of our proposed method, we conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [17], SpeFormer [18], MAEST [20], HiT [21], SSFTT [25], 3DViT [26], HUSST [27], and ViM [45]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references. The comparison is grounded in the same number of the training samples.
@@
    SSFTT combines 3D-CNNs and 2D-CNNs, featuring a Gaussian-weighted feature tokenizer to prepare extracted features for the Transformer encoder. In our experiments, patch sizes vary by dataset: $13\times 13$ for the IP and PU dataset and $9\times 9$ for both HU2013 dataset.
@@
    Our MiM model standardizes all feature dimensions at 64, with T-Mamba encoders having a depth of 2. The initial patch sizes are set at $7\times 7$ for the IP dataset, $11\times 1$ 11 for the PU dataset, and $9\times 9$ for the HU2013 dataset. Other settings can be found in Algorithm 1.
@@ IV-D1 Quantitative Results and Classification Maps
Tables II, III, and IV present detailed classification metrics such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, and HU2013 datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, and 12.
@@
TABLE V: TOTAL RUNNING TIME (MINUTES) AND THE NUMBER OF PARAMETERS REQUIRED BY DIFFERENT METHODS ON THE INDIAN PINES DATASET.
@@
Parameters 416K     644K     953K     45.8M     2.9M     2.4M     9.1M     357K     1.7M
Our method, MiM, introduces a MCS mechanism with a T-Mamba encoder, facilitating lightweight, multi-directional feature generation. Compared to ViM, our specialized T-Mamba encoder and extended bidirectional scan improve classification performance significantly. Our approach, notably lighter than HiT, SSFTT, and HUSST, produces classification maps with smoother regions and more distinct land-cover boundaries. The class-wise accuracies are competitive and generally superior to other methods. Key differences in classification performance are highlighted and marked in Figs. 10, 11, and 12 for detailed comparisons.
@@ IV-D3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table V. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset. All tests were run on an Ubuntu 22.04 system equipped with a GeForce RTX 3090 GPU.
@@ IV-E Other Analysis and Ablation Study
TABLE VI: IMPACT OF DIFFERENT INSERTABLE COMPONENTS IN T-MAMBA ENCODER ON THE CLASSIFICATION RESULTS (%). THE BEST ONE IS HIGHLIGHTED AS RED.
$STL()$     $GDM()$     $STF()$ IP     PU     HU2013
 1 ✗     ✗     82.559     80.999     88.112
2     ✗     86.129     85.698     90.995
3 ✗         84.006     82.159     89.459
4 ✗     ✗     84.561     83.226     90.642
5         87.021     86.111     91.324
6     ✗     91.969     89.123     91.452
7 ✗         87.159     87.661     91.003
8         92.079     91.576     92.892
@@ IV-E1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table VI. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
@@
TABLE VII: IMPACT OF DIFFERENT INITIAL PATCH SIZE FOR THE OVERALL ACCURACY (%) ON THREE DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
Initial Patch Size ($p$) IP     PU     HU2013
 $3\times 3$ 89.001 $\uparrow$     86.331 $\uparrow$     89.123 $\uparrow$
$5\times 5$ 91.332 $\uparrow$     89.126 $\uparrow$     91.028 $\uparrow$
$7\times 7$ 92.079 $\uparrow$     90.214 $\uparrow$     $9$uparrow
$9\times 9$ 91.554 $\downarrow$     91.002 $\uparrow$     92.892 $\uparrow$
$11\times 1$ 11 90.667 $\downarrow$     91.576 $\uparrow$     92.091 $\downarrow$
$13\times 13$ 89.214 $\downarrow$     90.885 $\downarrow$     90.451 $\downarrow$
$15\times 15$ 87.332 $\downarrow$     88.213 $\downarrow$     88.519 $\downarrow$
@@ IV-E2 Analysis on Initial Patch Size
As shown in Table VII, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for three datasets, are detailed in the same table.
@@
For the relatively compact IP dataset, the ideal initial patch size is small, confirmed at 7 $\times$ 7. In contrast, for the HU2013 dataset, which features a dense spatial distribution but closely spaced land-covers, a slightly larger patch size of 9 $\times$ 9 is optimal. The PU dataset, characterized by its flat and smooth terrain, warrants an even larger optimal initial patch size of 11 $\times$ 11.
@@ IV-E3 Investigation of Scan Design
TABLE VIII: THE COMPARISON OF USING DIFFERENT SCAN DESIGNS FOR OA (%). THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
Num. of used types IP     PU     HU2013
  1 89.123     89.559     90.123
Mamba-scan 2 90.156     90.452     91.029
3 91.221     91.001     91.993
4 92.079     91.576     92.892
  1 88.326     87.009     89.163
Raster-scan 2 89.555     88.322     90.442
3 90.122     89.125     91.105
4 91.612     90.761     91.986
  1 88.039     86.002     89.013
Diagonal-scan 2 89.196     87.569     89.669
3 90.236     88.555     90.139
4 91.002     89.571     91.004
  1 88.129     87.563     87.125
Zig-Zag-scan 2 90.009     87.993     89.699
3 91.011     88.961     90.861
4 91.536     91.299     91.834
We tested these scanning methods to evaluate their performance, with results summarized in Table VIII. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
@@ IV-E4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 15: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across three datasets. The results of these experiments are detailed in Table IX.
TABLE IX: IMPACT OF DIFFERENT FEATURE FUSION STRATEGY IN T-MAMBA ENCODER FOR THE OA (%) ON THREE DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
IP     PU     HU2013
 Post-Merge 91.326     90.332     91.229
Pre-Merge 92.079     91.576     92.892
@@
Figure 16: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). Experimental results are listed in (a) IP dataset, (b) PU dataset, and (c) HU2013 dataset. Meanwhile, the (d) shows the training time, OA(%) and model size by using those networks.
@@ IV-E5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN [^8] https://pytorch.org/docs/stable/generated/torch.nn.RNN.html, GRU [^9] https://pytorch.org/docs/stable/generated/torch.nn.GRU.html, LSTM [^10] https://pytorch.org/docs/stable/generated/torch.nn.LSTM.html, general Transformer encoder [^11] https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder.html, RT (a hybrid of LSTM and Transformer) [38], and a general Mamba model [40]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. All tests were performed under identical conditions, with the results shown in Fig. 16.
@@
TABLE X: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA ON THE INDIAN PINES DATASET WITH DATA AUGMENTATION. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
IV-E6 Further Discussion on Mamba and Transformer Based under Data Augmentation
@@
The results, detailed in Table X, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.
However, when the training sample size was expanded six-fold to 4170, our MiM model was no longer the top performer. Instead, MAEST, which integrates the MAE with prior masks, achieved the best results. Transformer-based models generally performed well. This improvement is likely due to the Transformer’s mechanism, which effectively utilizes ample training samples to enhance its recognition and generalization capabilities, as supported by literature suggesting that Transformers can outperform CNNs when sufficient data is available.
Thus, while our MiM model excels in scenarios with limited, disjoint, and fixed samples—conditions that mimic real-world HSI applications, which would be beneficial for the few-shot learning. The Transformer-based methods tend to perform better under conditions of abundant training data.
@@ V Conclusion
In this study, we introduced a novel and pioneering approach to HSI classification utilizing the newly developed Mamba architecture. Specifically, we implemented a centralized MCS for multi-directional feature aggregation, creating a lighter and more streamlined framework for visual cognition alongside the Mamba sequential model. To further enhance this architecture, we designed a T-Mamba encoder, incorporating bi-directional GDM and a STL to refine feature quality and support downsampling-driven semantic token representation. Meanwhile, porposed STF is effectively fusing learned abstract representations with original sequence features. Additionally, we integrated a WMF module and a multi-scale loss function to optimize decoding efficiency. As the first application of a vision Mamba-based model for HSI classification, our results have significantly outperformed other baselines and state-of-the-art methods with limited and disjoint data samples. However, with augmented training samples, Transformer-based models tend to be better. Moving forward, we plan to implement our idea into the few-shot learning with small samples and extend this approach to include a 3D version of the Mamba model, expanding beyond the current 2D implementation.

### v3

@@ IV-C Brief Description and Settings of Compared Methods
To evaluate the effectiveness of our proposed method, we conducted comparative analyses against a range of baseline and state-of-the-art models. These included vanilla ViT (ViT) [18], SpeFormer [19], MAEST [21], HiT [22], SSFTT [26], 3DViT [27], HUSST [28], and ViM [52]. All models were employed as patch-wise classifiers, with parameter settings adhering to their respective references, especially the initial patch size for different datasets. The comparison is grounded in the same number of the training samples without data augmentation.
@@
    SSFTT combines 3D-CNNs and 2D-CNNs, featuring a Gaussian-weighted feature tokenizer to prepare extracted features for the Transformer encoder. In our experiments, patch sizes vary by dataset: $13\times 13$ for the IP and PU dataset and $9\times 9$ for both HU2013 and HongHu dataset.
@@
    Our MiM model standardizes all feature dimensions at 64, with T-Mamba encoders having a depth of 2. The initial patch sizes are set at $7\times 7$ for the IP dataset, $11\times 1$ 11 for the PU dataset, and $9\times 9$ for the HU2013 and HongHu dataset. Others can be found in Algorithm 1.
TABLE VI: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA, AA, AND Kappa AS WELL AS THE ACCURACIES FOR EACH CLASS ON THE WHU-Hi-HongHu DATASET. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
 1         83.308     92.414     93.150     92.397     94.062     93.913     93.392     90.774     91.922
2         51.588     77.377     69.992     83.118     80.321     84.327     82.783     78.092     83.239
3         66.129     87.049     87.753     86.075     86.447     88.706     86.066     85.647     89.772
4         81.673     92.768     95.519     95.807     92.688     93.738     94.840     90.939     95.993
5         30.789     42.216     52.979     50.280     39.060     45.475     59.836     40.896     59.971
6         80.313     93.287     94.380     94.295     93.865     93.009     94.936     92.830     96.003
7         55.159     69.221     81.213     66.949     84.649     74.366     79.445     81.871     83.873
8         29.825     48.326     45.361     47.745     46.911     64.771     45.650     54.258     64.170
9         87.712     95.937     95.455     93.624     96.133     94.108     95.912     95.570     96.399
10         50.244     78.034     85.437     78.864     82.318     85.436     83.367     77.348     87.094
11         32.714     68.521     80.407     64.692     76.228     82.055     72.825     75.274     88.261
12         38.737     60.166     72.607     64.964     76.981     65.600     69.342     68.752     77.643
13         41.373     66.599     75.962     60.533     72.911     70.993     75.059     66.550     76.611
14         53.612     75.796     88.200     85.221     85.478     90.432     86.554     74.109     89.186
15         48.922     80.839     84.730     83.587     84.911     70.855     72.665     37.689     87.631
16         85.969     96.110     94.660     94.505     95.649     95.244     96.007     93.233     95.948
17         71.222     89.171     74.695     94.118     85.529     87.956     87.207     85.676     92.016
18         49.649     84.597     83.587     85.643     83.333     84.548     86.010     89.097     86.875
19         59.138     83.529     86.927     82.922     90.572     86.385     88.418     86.205     91.872
20         34.296     71.906     80.832     32.043     79.922     80.556     82.181     74.823     88.583
21         20.711     50.048     52.946     52.932     46.456     62.122     44.932     56.133     71.989
22         32.600     66.921     80.489     87.111     68.578     83.324     67.961     53.190     83.057
 OA         65.2688     82.9934     87.3104     84.2750     85.2845     85.8091     86.9473     82.7789     89.5749
AA         67.4351     84.6977     85.7987     83.7510     87.4042     88.7126     87.9006     85.7671     90.9751
Kappa         0.5873     0.7911     0.8427     0.8048     0.8196     0.8254     0.8387     0.7897     0.8705
Figure 13: The classification maps with different methods for the HongHu dataset. (a) vanilla ViT (OA=65.2688%). (b) SpeFormer (OA=82.9934%). (c) MAEST (OA=87.3104%) (d) HiT (OA=84.2750%). (e) SSFTT (OA=85.2845%). (f) 3DViT (OA=85.8091%). (g) HUSST (OA=86.9473%). (h) ViM (OA=82.7789%). (i) MiM (OA=89.5749%).
@@ IV-D1 Quantitative Results and Classification Maps
Tables III, IV, V, and VI present average classification metrics under 5 times of running, such as OA, AA, Kappa coefficient, and class-wise accuracy for the IP, PU, HU2013, and HongHu datasets, respectively. Corresponding classification maps are displayed in Figs. 10, 11, 12, and 13.
@@
TABLE VII: TOTAL RUNNING TIME (MINUTES) AND THE NUMBER OF PARAMETERS REQUIRED BY DIFFERENT METHODS ON THE INDIAN PINES DATASET.
@@
Parameters 416K     644K     953K     45.8M     2.9M     2.4M     9.1M     357K     1.1M
Our method, MiM, introduces a MCS mechanism with a T-Mamba encoder, facilitating lightweight, multi-directional feature generation. Compared to ViM, our specialized T-Mamba encoder and extended bidirectional scan improve classification performance significantly. Our approach, notably lighter than HiT, SSFTT, and HUSST, produces classification maps with smoother regions and more distinct land-cover boundaries. The class-wise accuracies are competitive and generally superior to other methods.
@@ IV-D3 Comparison of Computational Costs
To evaluate computational efficiency across all methods, we conducted a detailed statistical analysis, summarized in Table VII. This analysis adheres to previously described experimental setups and includes metrics such as total training time (minutes), testing time (seconds), and model size (number of parameters) for the IP dataset.
@@ IV-E Other Analysis and Ablation Study
TABLE VIII: IMPACT OF DIFFERENT INSERTABLE COMPONENTS IN T-MAMBA ENCODER ON THE CLASSIFICATION RESULTS (%). THE BEST ONE IS HIGHLIGHTED AS RED.
$STL$     $GDM$     $STF$ IP     PU     HU2013     HongHu
 ✗     ✗     82.559     80.999     88.112     84.123
    ✗     86.129     85.698     90.995     86.889
✗         84.006     82.159     89.459     86.114
✗     ✗     84.561     83.226     90.642     85.013
        87.021     86.111     91.324     88.546
    ✗     91.969     89.123     91.452     88.982
✗         87.159     87.661     91.003     88.002
        92.079     91.576     92.892     89.575
@@ IV-E1 Ablation on Proposed Different Components
In the T-Mamba encoder, we integrated several key components: the STL, GDM, and STF. To assess their effectiveness, we conducted an ablation study, the results of which are summarized in Table VIII. The ablation study details how each component was tested. In scenarios where STL and STF were not used, we substituted them with general average pooling and simple summation, respectively, to gauge their relative performance.
@@
TABLE IX: IMPACT OF DIFFERENT INITIAL PATCH SIZE FOR THE OVERALL ACCURACY (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
Patch Size ($p$) IP     PU     HU2013     HongHu
 $3\times 3$ 89.001 $\uparrow$     86.331 $\uparrow$     89.123 $\uparrow$     83.127 $\uparrow$
$5\times 5$ 91.332 $\uparrow$     89.126 $\uparrow$     91.028 $\uparrow$     87.269 $\uparrow$
$7\times 7$ 92.079 $\uparrow$     90.214 $\uparrow$     $9$uparrow     89.113 $\uparrow$
$9\times 9$ 91.554 $\downarrow$     91.002 $\uparrow$     92.892 $\uparrow$     89.575 $\uparrow$
$11\times 1$ 11 90.667 $\downarrow$     91.576 $\uparrow$     92.091 $\downarrow$     88.990 $\downarrow$
$13\times 13$ 89.214 $\downarrow$     90.885 $\downarrow$     90.451 $\downarrow$     87.317 $\downarrow$
$15\times 15$ 87.332 $\downarrow$     88.213 $\downarrow$     88.519 $\downarrow$     86.451 $\downarrow$
@@ IV-E2 Analysis on Initial Patch Size
As shown in Table IX, we evaluated a range of initial patch sizes (3, 5, 7, 9, 11, 13, 15). The results, specifically OA for four datasets, are detailed in the same table.
@@
For the relatively compact IP dataset, the ideal initial patch size is small, confirmed at 7 $\times$ 7. In contrast, for the HU2013 dataset, which features a dense spatial distribution but closely spaced land-covers, a slightly larger patch size of 9 $\times$ 9 is optimal. The PU dataset, characterized by its flat and smooth terrain, warrants an even larger optimal initial patch size of 11 $\times$ 11. Finally, the HongHu dataset gets the ideal initial patch size at 9 $\times$ 9.
@@ IV-E3 Investigation of Scan Design
TABLE X: THE COMPARISON OF USING DIFFERENT SCAN DESIGNS FOR OA (%). THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
Num. of used IP     PU     HU2013     HongHu
  1 89.123     89.559     90.123     87.315
Mamba-scan 2 90.156     90.452     91.029     88.569
3 91.221     91.001     91.993     89.226
4 92.079     91.576     92.892     89.575
  1 88.326     87.009     89.163     85.429
Raster-scan 2 89.555     88.322     90.442     86.129
3 90.122     89.125     91.105     88.226
4 91.612     90.761     91.986     89.125
  1 88.039     86.002     89.013     83.664
Diagonal-scan 2 89.196     87.569     89.669     86.548
3 90.236     88.555     90.139     87.841
4 91.002     89.571     91.004     88.221
  1 88.129     87.563     87.125     84.003
Zig-Zag-scan 2 90.009     87.993     89.699     86.443
3 91.011     88.961     90.861     87.231
4 91.536     91.299     91.834     87.992
We tested these scanning methods to evaluate their performance, with results summarized in Table X. Each scan design was tested using one to four types. The most effective results were achieved using the Mamba-scan with four directional types, confirming that a continuous sequence arrangement enhances content recognition. This finding is further supported by comparing the Diagonal-scan and the Zig-zag-scan, where the latter showed slight improvements due to its semi-continuous positioning. The commonly used Raster-scan, despite its prevalence in image transformation tasks, tends to disrupt spatial continuity, particularly when spatial context is significant in RS land-cover analysis, resulting in inferior performance compared to the Mamba-scan.
@@ IV-E4 Exploring the Feature Fusion Strategy in T-Mamba
We developed our bi-directional feature fusion strategy in T-Mamba encoder with two ways, as shown in Fig. 16: 1) Post-Merge strategy, and 2) Pre-Merge strategy. To evaluate the effectiveness of these two strategies, we conducted experiments across four datasets. The results of these experiments are detailed in Table XI.
TABLE XI: IMPACT OF DIFFERENT FEATURE FUSION STRATEGY IN T-MAMBA ENCODER FOR THE OA (%) ON FOUR DATASETS. THE BEST ONE IS HIGHLIGHTED AS RED.
IP     PU     HU2013     HongHu
 Post-Merge 91.326     90.332     91.229     88.450
Pre-Merge 92.079     91.576     92.892     89.575
@@
Figure 17: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). We have to note that all networks are not sharing parameters for four types of MCS in MiM model. Experimental results are listed in (a) IP dataset, (b) PU dataset, (c) HU2013 dataset, and (d) HongHu dataset. Meanwhile, the (e) shows the training time, OA(%) and model size by using those networks.
@@ IV-E5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN [^10] https://pytorch.org/docs/stable/generated/torch.nn.RNN.html, GRU [^11] https://pytorch.org/docs/stable/generated/torch.nn.GRU.html, LSTM [^12] https://pytorch.org/docs/stable/generated/torch.nn.LSTM.html, general Transformer encoder [^13] https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder.html, RT (a hybrid of LSTM and Transformer) [45], and a general Mamba model [47]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. We have to note that all networks are not sharing parameters for four types of MCS in MiM model. All tests were performed under identical conditions, with the results shown in Fig. 17.
@@
TABLE XII: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA ON THE INDIAN PINES DATASET WITH DATA AUGMENTATION. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
@@
IV-E6 Further Discussion on Mamba and Transformer under Data Augmentation
@@
The results, detailed in Table XII, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.
However, when the number of training samples are expanded six-fold to 4170, our MiM model was no longer the top performer. Instead, MAEST, which integrates the MAE with prior masks, achieved the best results. Transformer-based models generally performed well. This improvement is likely due to the Transformer’s mechanism, which effectively utilizes ample training samples to enhance its recognition and generalization capabilities, as supported by literature suggesting that Transformers can outperform CNNs when sufficient data is available.
Thus, while our MiM model excels in scenarios with limited, disjoint, and fixed samples—conditions that mimic real-world HSI applications. The Transformer-based methods tend to perform better under conditions of abundant training data.
@@ V Conclusion
In this study, we introduced a novel and pioneering approach to HSI classification utilizing the newly developed Mamba architecture. Specifically, we implemented a centralized MCS for multi-directional feature aggregation, creating a lighter and more streamlined framework for visual cognition alongside the Mamba sequential model. To further enhance this architecture, we designed a T-Mamba encoder, incorporating bi-directional GDM and a STL to refine feature quality and support downsampling-driven semantic token representation. Meanwhile, porposed STF is effectively fusing learned abstract representations with original sequence features. Additionally, we integrated a WMF module and a multi-scale loss function to optimize decoding efficiency. As the pioneering application of a vision Mamba-based model for HSI classification, our results have significantly outperformed other baselines and state-of-the-art methods with limited and disjoint data samples. However, with augmented training samples, Transformer-based models tend to be better. Moving forward, we plan to extend our idea to include a 3D version of the Mamba model where the scanning consistency on spectral information may work.